Constants Periodic Table PartA n electron with speed 2 55* 10 m,/s straveling paralel to a uniform electric field d magnitude 1.18 104 N/C How far will the electron travel before it stops? In Submit Previous Answers Request Answer X Incorrect; Try Again: 4 attempts remaining Part B How much time will elapse before it returns to its starting point? Provide Feedback
Answer #1

Either straight-line kinematics or Work-Energy equations should give the right answer.

The decelerating force is F = qE = -1.6*10^-19 * 1.18*10^4 = -1.888*10^-15 N
so the deceleration is a = F/m = -1.888*10^-15 / 9.11*10^-31kg = -2.0724*10^15 m/s^2

a)
Vf^2=0 = 2ad+Vo^2
so d = -Vo^2 / 2a = -(2.55*10^7)^2 / -2*2.0724*10^15 = 0.1568 mm

d = 1.586*10^-4m

b)
time to stop is t = -Vo/a = -2.55*10^7 / -2.0724*10^15 = 1.230*10^-8 s